Osmotic Pressure question: Looking for the molar weight of protein with the following numbers:

Started with a sample of 1.928 grams of protein

30 degrees Celcius 66.3 mATM

40 mL of water

what is the molar weight of protein in g/mol???

*Gives a hint that the weight should be 10^4 - 10^6 g/mol
Answer #1

P = MRT

for

m = 1.928 g of protein

T= 30°C = 303 K

P = 66.3*10^-3 atm

V = 40*10^-3 L

so

substitute

M = P/(RT) = (66.3*10^-3) / (0.082*303) = 0.0026684 mol per liter

mol = MV = 0.0026684 * 40*10^-3 = 0.000106736 mol

MW = mass/mol = (1.928) / (0.000106736 = 18063.2588 which is in the reange of 10^4